# Dynamical symmetry breaking in Georgi-Glashow chiral-gauge theories

### v1

@@ V Conclusions
We constructed the Fierz-complete basis of four-fermion operators and derived their flows using the functional renormalization group. These carry the information on the resonant structure of the different channels and easily permit to compute the critical coupling necessary to trigger the dynamical generation of a scale. We find that the phase structure of generalized GG theories is rich as various regions with distinct patterns of dSB. Such richness is a consequence of considering a theory with different fermion species rather than from the inherently chiral nature, as this provides strong constraints on the inducible operators. We note that, to the best of our knowledge, such type of multi-species gauge-fermion theories have not been studied before with approaches sensitive to the off-shell dynamics.
@@
The present analysis opens the door to many future studies. In an upcoming companion paper, we will extend this work in the symmetric case, commonly known as the Bars-Yankielowicz Bars and Yankielowicz (1981) model, exploring whether similar or qualitatively new symmetry-breaking patterns arise. Furthermore, the IR dynamics of these models can be tackled from first-principles in the present framework by dynamical bosonizing the dominant channel and including the dynamical appearance of a gauge-boson mass gap. More broadly, the functional method can be applied to investigate the mechanism of symmetric mass generation in various spacetime dimensions, see e.g. Wang and You (2022) for a review. Once all the ’t Hooft anomalies are canceled, this mechanism can be used to generate a mass gap for chiral fermions without breaking the chiral symmetry. See e.g. Wang and Wen (2023, 2018); Zeng et al. (2022); Tong (2022) for the studies on the chiral fermion 3 - 4 - 5 - 0 model in 1+1 dimensions, and Razamat and Tong (2021); Tong (2022) for other related models in 3+1 dimensions.
@@ Acknowledgements.
We would like to thank Jan M. Pawlowski for helpful comments and suggestions. APG is supported by the RIKEN Special Postdoctoral Researcher (SPDR) Program. SV is supported by the Fonds de la Recherche Scientifique de Belgique (FNRS) under the IISN convention 4.4517.08. We are grateful to the Mainz Institute for Theoretical Physics (MITP) of the DFG Cluster of Excellence PRISMA+ (Project ID 39083149) for its hospitality and support during the initial stages of this work. H-.L.L. is supported by the start-up funding of Sun Yat-Sen University under grant number 74130-12255013. The work of L.X.X. is partially supported by “Exotic High Energy Phenomenology” (X-HEP), a project funded by the European Union - Grant Agreement n.101039756. This project is also supported by the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP) which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311.
@@ Appendix A Four-fermion operator construction
This Appendix is devoted to the systematic construction of the four-fermion operators in generalized GG theories. We demand that operators are not only invariant under the $SU(N_{c})$ gauge group, but also under the global flavor symmetry for each chiral fermion and the residual anomaly-free $U(1)$ symmetry as well. This allows only three operator types: ${\chi^{\dagger}}^{2}\chi^{2}$, ${\chi^{\dagger}}\chi{\psi^{\dagger}}\psi$ and ${\psi^{\dagger}}^{2}\psi^{2}$, except for the scenario $N_{c}=5$ with $N_{\textrm{gen}}=1$ in the GG model, where one additional operator type (and its complex conjugate) is allowed: $\psi\chi^{3}$ (see discussion in III.
@@
with $f_{1,2}$ the flavor indices for the field $\psi$ and $f_{3,4}$ for the field ${\psi^{\dagger}}$, and the subscript $S,A$ indicates that the tensor is symmetric or anti-symmetric under the permutation of 1,21,21 , 2 or 3,43,43 , 4. Similarly, the gauge tensor can also be classified as $T^{[g]}_{S}$ and $T^{[g]}_{A}$ as follows:
@@
where $T_{f}$ is the generator of fundamental representation. The advantage to use this notation is that in the computation of the flow equation, the trace involving the antisymmetric generators can always be converted to the trace of the fundamental generators, which can be simplified with programs such as FeynCalc Shtabovenko et al. (2025).
@@ Appendix B Full effective action and flow equations
[…]
@@
The most efficient strategy is to construct projection operators such that each projected correlator depends solely on one of the dressing functions Braun et al. (2025); Gehring et al. (2015). This is in fact, particularly relevant when bosonizing the theory. A simpler alternative, sufficient for the present purpose, is to construct a complete basis of projectors directly adapted to the tensor structures under consideration. We define:
@@
$\partial_{t}\bar{\lambda}_{1}$ $=(2+2\eta_{\psi})\bar{\lambda}_{1}+\frac{k^{2}}{Z_{\psi}^{2}}\frac{\partial_{t}\left[{\cal P}_{1,R}^{abcd}\Gamma_{k}^{(\bar{\psi}_{a}\psi_{b}\bar{\psi}_{c}\psi_{d})}\right](N_{c}N_{\psi}+1)-\partial_{t}\left[{\cal P}_{2,R}^{abcd}\Gamma_{k}^{(\bar{\psi}_{a}\psi_{b}\bar{\psi}_{c}\psi_{d})}\right](N_{c}+N_{\psi})}{128N_{c}(N_{c}^{2}-1)N_{\psi}(N_{\psi}^{2}-1)}\,,$ (26a)
$\partial_{t}\bar{\lambda}_{2}$ $=(2+2\eta_{\psi})\bar{\lambda}_{2}+\frac{k^{2}}{Z_{\psi}^{2}}\frac{\partial_{t}\left[{\cal P}_{2,R}^{abcd}\Gamma_{k}^{(\bar{\psi}_{a}\psi_{b}\bar{\psi}_{c}\psi_{d})}\right](N_{c}N_{\psi}+1)-\partial_{t}\left[{\cal P}_{1,R}^{abcd}\Gamma_{k}^{(\bar{\psi}_{a}\psi_{b}\bar{\psi}_{c}\psi_{d})}\right](N_{c}+N_{\psi})}{128N_{c}(N_{c}^{2}-1)N_{\psi}(N_{\psi}^{2}-1)}\,.$ (26b)
@@
$\partial_{t}\bar{\lambda}_{1}=$ $(2+2\,\eta_{\psi})\bar{\lambda}_{1}+\frac{1}{16\pi^{2}}\bigg{[}\frac{g^{4}\left(3N_{c}^{2}+4\right)(5\eta_{A}+3\eta_{\psi}-45)}{160N_{c}^{2}}+\frac{g^{2}(5\eta_{A}+6\eta_{\psi}-60)(\bar{\lambda}_{1}-\bar{\lambda}_{2}N_{c})}{10N_{c}}$ (27a)
$-\frac{1}{20}(\eta_{\chi}-5)N_{\chi}\left(\bar{\lambda}_{6}^{2}(N_{c}-1)N_{c}+\frac{(N_{c}-2)}{2N_{c}}\bar{\lambda}_{7}^{2}\right)+\frac{4}{5}(\eta_{\psi}-5)\left(\bar{\lambda}_{2}^{2}-\bar{\lambda}_{1}\bar{\lambda}_{2}(N_{c}+N_{\psi})-\bar{\lambda}_{1}^{2}\frac{(N_{c}N_{\psi}-1)}{2}\right)\bigg{]}\,,$
$\partial_{t}\bar{\lambda}_{2}=$ $(2+2\,\eta_{\psi})\bar{\lambda}_{2}+\frac{1}{16\pi^{2}}\bigg{[}\frac{g^{4}\left(N_{c}^{2}-8\right)(5\eta_{A}+3\eta_{\psi}-45)}{160N_{c}}+\frac{g^{2}(5\eta_{A}+6\eta_{\psi}-60)(\bar{\lambda}_{2}-\bar{\lambda}_{1}N_{c})}{10N_{c}}$
@@
$\partial_{t}\bar{\lambda}_{3}=$ $(2+2\,\eta_{\chi})\bar{\lambda}_{3}+\frac{1}{16\pi^{2}}\bigg{[}\frac{1}{10}g^{4}(5\eta_{A}+3\eta_{\chi}-45)+\frac{\left(g^{2}(5\eta_{A}+6\eta_{\chi}-60)\left(N_{c}\bar{\lambda}_{5}-2\bar{\lambda}_{3}\right)\right)}{5N_{c}}$
$+\frac{1}{5}(\eta_{\chi}-5)\left(8\bar{\lambda}_{4}\bar{\lambda}_{3}+\frac{4\left(N_{c}^{2}-N_{c}+4\right)}{N_{c}}\bar{\lambda}_{5}\bar{\lambda}_{3}-\bar{\lambda}_{3}^{2}\left(N_{c}^{2}-N_{c}+2N_{\chi}\right)+6\bar{\lambda}_{5}^{2}\right)\bigg{]}\,,$ (27c)
$\partial_{t}\bar{\lambda}_{4}=$ $(2+2\,\eta_{\chi})\bar{\lambda}_{4}+\frac{1}{16\pi^{2}}\bigg{[}\frac{g^{4}\left(N_{c}^{2}-2N_{c}-4\right)(5\eta_{A}+3\eta_{\chi}-45)}{10N_{c}^{2}}$
$-g^{2}\frac{\left(5\eta_{A}+6\eta_{\chi}-60\right)\left(\lambda_{5}\left(N^{2}{}_{c}-2N_{c}-4\right)+2\lambda_{3}N_{c}\right)}{5N_{c}^{2}}$ (27d)
@@
$+\frac{\left(\eta_{\chi}-5\right)}{10}\left(-4\bar{\lambda}_{4}\bar{\lambda}_{3}\left(N_{c}^{2}-N_{c}+2N_{\chi}\right)+\bar{\lambda}_{4}^{2}\left(-2N_{\chi}N_{c}^{2}+2N_{\chi}N_{c}+4\right)+8\bar{\lambda}_{3}^{2}\right)-\frac{\left(\eta_{\psi}-5\right)}{10}N_{\psi}N_{c}\bar{\lambda}_{6}^{2}\bigg{]}\,,$ (27e)
$\partial_{t}\bar{\lambda}_{5}=$ $(2+2\,\eta_{\chi})\bar{\lambda}_{5}+\frac{1}{16\pi^{2}}\bigg{[}\frac{g^{4}\left(N_{c}^{2}-8N_{c}-32\right)(5\eta_{A}+3\eta_{\chi}-45)}{80N_{c}}+\frac{g^{2}(5\eta_{A}+6\eta_{\chi}-60)(\bar{\lambda}_{5}(N_{c}+4)-N_{c}(\bar{\lambda}_{4}+\bar{\lambda}_{3}))}{5N_{c}}$
$+\frac{1}{5}(\eta_{\chi}-5)\left(8\bar{\lambda}_{4}\bar{\lambda}_{5}-4\bar{\lambda}_{3}\bar{\lambda}_{5}(N_{\chi}-3)-\frac{\left(N_{c}^{2}(N_{\chi}+3)-2N_{c}(N_{\chi}-1)+16\right)}{N_{c}}\bar{\lambda}_{5}^{2}\right)-\frac{1}{20}(\eta_{\psi}-5)\bar{\lambda}_{7}^{2}N_{\psi}\bigg{]}\,,$ (27f)
$\partial_{t}\bar{\lambda}_{6}=$ $(2+\,\eta_{\psi}+\,\eta_{\chi})\bar{\lambda}_{6}+\frac{1}{16\pi^{2}}\bigg{[}\frac{g^{4}\left(N_{c}^{2}-N_{c}-2\right)(10\eta_{A}+3(\eta_{\chi}+\eta_{\psi}-30))}{20N_{c}^{2}}$
$-\frac{g^{2}\bar{\lambda}_{7}\left(N_{c}^{2}-N_{c}-2\right)(5\eta_{A}+3(\eta_{\chi}+\eta_{\psi}-20))}{10N_{c}^{2}}+\frac{3(\eta_{\chi}+\eta_{\psi}-10)}{10}\left(\bar{\lambda}_{6}^{2}+\frac{\left(N_{c}^{2}-N_{c}-2\right)}{2N_{c}^{2}}\bar{\lambda}_{7}^{2}\right)$
$-\frac{1}{5}(\eta_{\chi}-5)\left(\bar{\lambda}_{4}\bar{\lambda}_{6}\left(N_{c}^{2}N_{\chi}-N_{c}N_{\chi}+2\right)+\bar{\lambda}_{3}\bar{\lambda}_{6}\left(N_{c}^{2}-N_{c}+2N_{\chi}\right)+\frac{2\left(N_{c}^{2}-N_{c}-2\right)}{N_{c}}\bar{\lambda}_{5}\bar{\lambda}_{6}\right)$
$-\frac{2}{5}(\eta_{\psi}-5)\left(\bar{\lambda}_{1}\bar{\lambda}_{6}(N_{c}N_{\psi}+1)+\bar{\lambda}_{2}\bar{\lambda}_{6}(N_{c}+N_{\psi})\right)\bigg{]}\,,$ (27g)
$\partial_{t}\bar{\lambda}_{7}=$ $(2+\,\eta_{\psi}+\,\eta_{\chi})\bar{\lambda}_{7}+\frac{1}{16\pi^{2}}\bigg{[}\frac{g^{4}\left(N_{c}^{2}-4N_{c}-16\right)(10\eta_{A}+3(\eta_{\chi}+\eta_{\psi}-30))}{80N_{c}}$
@@
$+\frac{(\eta_{\chi}+\eta_{\psi}-10)}{5}\left(3\bar{\lambda}_{6}\bar{\lambda}_{7}-\frac{\left(N_{c}^{2}+3N_{c}+12\right)}{4N_{c}}\bar{\lambda}_{7}^{2}\right)-\frac{2(\eta_{\psi}-5)}{5}\left(\bar{\lambda}_{1}\bar{\lambda}_{7}+\bar{\lambda}_{2}\bar{\lambda}_{7}N_{\psi}\right)$
$-\frac{2(\eta_{\chi}-5)}{5}\left(\bar{\lambda}_{4}\bar{\lambda}_{7}+\bar{\lambda}_{3}\bar{\lambda}_{7}N_{\chi}-\frac{\left(N_{c}^{2}(N_{\chi}+1)-2N_{c}(N_{\chi}+1)-4\right)}{10N_{c}}\bar{\lambda}_{5}\bar{\lambda}_{7}\right)\bigg{]}\,.$ (27h)
@@
$\eta_{A}=$ $-\frac{\partial_{t}Z_{A}}{Z_{A}}=-\left.\frac{\partial_{p^{2}}\,{\rm tr}\left[{\cal P}^{(AA)}\partial_{t}\Gamma^{(AA)}_{k}\right]}{Z_{A}\,{\rm tr}\left[{\left({\cal P}^{(AA)}\right)}^{2}\right]}\right|_{p=0}=\frac{g^{2}}{16\pi^{2}}\left[\left(\frac{31}{36}\eta_{A}-\frac{11}{3}\right)N_{c}-\frac{(\eta_{\chi}-4)}{6}(N_{c}-2)N_{\chi}-\frac{(\eta_{\psi}-4)}{6}N_{\psi}\right]\,,$ (28)
@@
and here we resum them, which provides one of the sources for the non-perturbative character of this approach. The $\cal P^{(\ldots)}$ stand for the projectors onto the kinetic terms of the respective fields.
Furthermore, from the flows in (27), the previously discussed coefficients ($\boldsymbol{c}_{\ldots}$) can be read off as a function of the number of chiral matter. To facilitate the identification of each contribution, we have displayed the flows as a function of the fermion multiplicities given by the global symmetries,
@@
So far, we have not specified the pure gauge and Dirac parts of the action. In this work we employ a truncation in this sectors which includes the interaction terms up to dimension four and the classical tensor structures which is the same one as in Goertz et al. (2024). Furthermore, in the gauge fixing sector we have used the Landau gauge. As an additional cross-check, we have derived the flow of the gauge couplings from all the included avatars and obtained the correct one-loop structure.
@@
Figure 2: Value of the $\boldsymbol{c}_{{\rm A},i}$ coefficients which reflect the strength of the primer gauge corrections on the generation of the different four-fermion dressings. The sign crossing of $\boldsymbol{c}_{{\rm A},5}$ at $N_{c}\sim 11$ and $\boldsymbol{c}_{{\rm A},7}$ provides the identification of regimes and potential subdomains.
@@
Some further remarks are in order. First, note that only $\boldsymbol{c}_{{\rm C},ijk}$ depends on $N_{\textrm{gen}}$, because the only source of $N_{\textrm{gen}}$ dependence in the flow arises from pure fermionic loops, which contribute exclusively to $\bar{\lambda}_{i}\bar{\lambda}_{j}$ terms.
@@
Moreover, part of the system has a resemblance to QCD which permits to perform some cross checks. Operators ${\cal O}_{1}$ and ${\cal O}_{2}$ are equivalent to those known as (V-A) and (V-A)adj but here only with one chirality Gies and Jaeckel (2006); Braun (2012); Goertz et al. (2024). In fact, we find an agreement up to a factor 1/2 in the $N_{\psi}$ contributions in their respective flows, providing a consistency check to our computation.

### v2

@@ 5 Conclusions
We constructed the Fierz-complete basis of four-fermion operators and derived their flows using the functional renormalization group. These carry the information on the channels which allow for a resonant growth and easily permit to compute the critical coupling necessary to trigger the dynamical generation of a scale. We find that the phase structure of generalized GG theories is rich as various regions with distinct patterns of dSB exist. Such richness originates from the presence of multiple fermion species rather than from the chiral nature of the theory itself. This is as the latter imposes strong constraints on the allowed four-fermion operators, while the former introduces additional ones that mix different species. We note that, to the best of our knowledge, such type of multi-species gauge-fermion theories have not been studied before with approaches sensitive to the off-shell dynamics.
@@
The present analysis opens the door to many future studies. In an upcoming companion paper, we will extend this work in the symmetric case, commonly known as the Bars-Yankielowicz Bars:1981se model, exploring whether similar or qualitatively new symmetry-breaking patterns arise. Furthermore, the IR dynamics of these models can be tackled from first-principles in the present framework by dynamical bosonizing the dominant channel and including the dynamical appearance of a gauge-boson mass gap. More broadly, the functional method can be applied to investigate the mechanism of symmetric mass generation in various spacetime dimensions, see e.g. Wang:2022ucy for a review. Once all the ’t Hooft anomalies are canceled, this mechanism can be used to generate a mass gap for chiral fermions without breaking the chiral symmetry. See e.g. Wang:2013yta ; Wang:2018ugf ; Zeng:2022grc ; Tong:2021phe for the studies on the chiral fermion 3 - 4 - 5 - 0 model in 1+1 dimensions, and Razamat:2020kyf ; Tong:2021phe for other related models in 3+1 dimensions.
@@ Acknowledgements.
We would like to thank Jan M. Pawlowski for valuable comments and suggestions. APG is supported by the RIKEN Special Postdoctoral Researcher (SPDR) Program. SV is supported by the Fonds de la Recherche Scientifique de Belgique (FNRS) under the IISN convention 4.4517.08. We are grateful to the Mainz Institute for Theoretical Physics (MITP) of the DFG Cluster of Excellence PRISMA+ (Project ID 39083149) for its hospitality and support during the initial stages of this work. H-.L.L. is supported by the start-up funding of Sun Yat-Sen University under grant number 74130-12255013 and by the National Science Foundation of China under Grants No.1250050417. The work of L.X.X. is partially supported by “Exotic High Energy Phenomenology” (X-HEP), a project funded by the European Union - Grant Agreement n.101039756. This project is also supported by the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP) which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311.
@@ Appendix A Four-fermion operator construction
This Appendix is devoted to the systematic construction of the four-fermion operators in generalized GG theories. We demand that operators are not only invariant under the $SU(N_{c})$ gauge group, but also under the global flavor symmetry for each chiral fermion and the residual anomaly-free $U(1)$ symmetry as well. This allows only three operator types: ${\chi^{\dagger}}^{2}\chi^{2}$, ${\chi^{\dagger}}\chi{\psi^{\dagger}}\psi$ and ${\psi^{\dagger}}^{2}\psi^{2}$, except for the scenario $N_{c}=5$ with $N_{\textrm{gen}}=1$ in the GG model, where one additional operator type (and its complex conjugate) is allowed: $\psi\chi^{3}$ (see discussion in 3.
@@
with $f_{1,2}$ the flavor indices for the field $\psi$ and $f_{3,4}$ for the field $\psi^{\dagger}$}, and the subscript $S,A$ indicates that the tensor is symmetric or anti-symmetric under the permutation of $1,2$ or $3,4$. Similarly, the gauge tensor can also be classified as $T^{[g]}_{S}$ and $T^{[g]}_{A}$ as follows:
@@
where $T_{f}$ is the generator of the fundamental representation. The advantage of using this notation is that in the computation of the flow equation, the trace involving the antisymmetric generators can always be converted to the trace of the fundamental generators, which can be simplified with programs such as FeynCalc Shtabovenko:2023idz .
@@ Appendix B Full effective action and flow equations
We proceed providing the flows of the dressing functions for all four-fermion operators previously constructed. These are obtained directly from the diagrammatic flow of purely fermionic four-point functions, which can be systematically and consistently evaluated using automated symbolic tools Huber:2019dkb ; Pawlowski:2021tkk . The computation follows closely the derived in Gies:2005as ; Goertz:2024dnz , and we now review it in detail for the particular case of the pure-$\psi$ operators.
The diagrammatic flow of the four-$\psi$ vertex reads:
@@
The most efficient strategy is to construct projection operators such that each projected correlator depends solely on one of the dressing functions Braun:2025gvq ; Gehring:2015vja . This is in fact, particularly relevant when bosonizing the theory. A simpler alternative, sufficient for the present purpose, is to construct a complete basis of projectors directly adapted to the tensor structures under consideration. We define:
@@
$\partial_{t}\bar{\lambda}_{1}$ $=(2+2\eta_{\psi})\bar{\lambda}_{1}$ (26a)
$\;\;+\frac{k^{2}}{Z_{\psi}^{2}}\frac{\partial_{t}\left[{\cal P}_{1,R}^{abcd}\Gamma_{k}^{(\bar{\psi}_{a}\psi_{b}\bar{\psi}_{c}\psi_{d})}\right](N_{c}N_{\psi}+1)-\partial_{t}\left[{\cal P}_{2,R}^{abcd}\Gamma_{k}^{(\bar{\psi}_{a}\psi_{b}\bar{\psi}_{c}\psi_{d})}\right](N_{c}+N_{\psi})}{128N_{c}(N_{c}^{2}-1)N_{\psi}(N_{\psi}^{2}-1)}\,$,
$\partial_{t}\bar{\lambda}_{2}$ $=(2+2\eta_{\psi})\bar{\lambda}_{2}$ (26b)
$\;\;+\frac{k^{2}}{Z_{\psi}^{2}}\frac{\partial_{t}\left[{\cal P}_{2,R}^{abcd}\Gamma_{k}^{(\bar{\psi}_{a}\psi_{b}\bar{\psi}_{c}\psi_{d})}\right](N_{c}N_{\psi}+1)-\partial_{t}\left[{\cal P}_{1,R}^{abcd}\Gamma_{k}^{(\bar{\psi}_{a}\psi_{b}\bar{\psi}_{c}\psi_{d})}\right](N_{c}+N_{\psi})}{128N_{c}(N_{c}^{2}-1)N_{\psi}(N_{\psi}^{2}-1)}\,.$
@@
$\partial_{t}\bar{\lambda}_{1}=$ $(2+2\,\eta_{\psi})\bar{\lambda}_{1}$
$+\frac{1}{16\pi^{2}}\bigg[\frac{g^{4}\left(3N_{c}^{2}+4\right)(5\eta_{A}+3\eta_{\psi}-45)}{160N_{c}^{2}}+\frac{g^{2}(5\eta_{A}+6\eta_{\psi}-60)(\bar{\lambda}_{1}-\bar{\lambda}_{2}N_{c})}{10N_{c}}$
$-\frac{1}{20}(\eta_{\chi}-5)N_{\chi}\left(\bar{\lambda}_{6}^{2}(N_{c}-1)N_{c}+\frac{(N_{c}-2)}{2N_{c}}\bar{\lambda}_{7}^{2}\right)$
$+\frac{4}{5}(\eta_{\psi}-5)\left(\bar{\lambda}_{2}^{2}-\bar{\lambda}_{1}\bar{\lambda}_{2}(N_{c}+N_{\psi})-\bar{\lambda}_{1}^{2}\frac{(N_{c}N_{\psi}-1)}{2}\right)\bigg]\,$, (27a)
$\partial_{t}\bar{\lambda}_{2}=$ $(2+2\,\eta_{\psi})\bar{\lambda}_{2}$
$+\frac{1}{16\pi^{2}}\bigg[\frac{g^{4}\left(N_{c}^{2}-8\right)(5\eta_{A}+3\eta_{\psi}-45)}{160N_{c}}+\frac{g^{2}(5\eta_{A}+6\eta_{\psi}-60)(\bar{\lambda}_{2}-\bar{\lambda}_{1}N_{c})}{10N_{c}}$
@@
$\partial_{t}\bar{\lambda}_{3}=$ $(2+2\,\eta_{\chi})\bar{\lambda}_{3}$
$+\frac{1}{16\pi^{2}}\bigg[\frac{1}{10}g^{4}(5\eta_{A}+3\eta_{\chi}-45)+\frac{\left(g^{2}(5\eta_{A}+6\eta_{\chi}-60)\left(N_{c}\bar{\lambda}_{5}-2\bar{\lambda}_{3}\right)\right)}{5N_{c}}$
$+\frac{\eta_{\chi}-5}{5}\left(8\bar{\lambda}_{4}\bar{\lambda}_{3}+\frac{4\left(N_{c}^{2}-N_{c}+4\right)}{N_{c}}\bar{\lambda}_{5}\bar{\lambda}_{3}-\bar{\lambda}_{3}^{2}\left(N_{c}^{2}-N_{c}+2N_{\chi}\right)+6\bar{\lambda}_{5}^{2}\right)\bigg]\,$, (27c)
$\partial_{t}\bar{\lambda}_{4}=$ $(2+2\,\eta_{\chi})\bar{\lambda}_{4}$
$+\frac{1}{16\pi^{2}}\bigg[\frac{g^{4}\left(N_{c}^{2}-2N_{c}-4\right)(5\eta_{A}+3\eta_{\chi}-45)}{10N_{c}^{2}}$
$-g^{2}\frac{\left(5\eta_{A}+6\eta_{\chi}-60\right)\left(\lambda_{5}\left(N^{2}{}_{c}-2N_{c}-4\right)+2\lambda_{3}N_{c}\right)}{5N_{c}^{2}}$
@@
$+\frac{\left(\eta_{\chi}-5\right)}{10}\left(-4\bar{\lambda}_{4}\bar{\lambda}_{3}\left(N_{c}^{2}-N_{c}+2N_{\chi}\right)+\bar{\lambda}_{4}^{2}\left(-2N_{\chi}N_{c}^{2}+2N_{\chi}N_{c}+4\right)+8\bar{\lambda}_{3}^{2}\right)$
$-\frac{\left(\eta_{\psi}-5\right)}{10}N_{\psi}N_{c}\bar{\lambda}_{6}^{2}\bigg]\,$, (27d)
$\partial_{t}\bar{\lambda}_{5}=$ $(2+2\,\eta_{\chi})\bar{\lambda}_{5}$
$+\frac{1}{16\pi^{2}}\bigg[\frac{g^{4}\left(N_{c}^{2}-8N_{c}-32\right)(5\eta_{A}+3\eta_{\chi}-45)}{80N_{c}}$
$+\frac{g^{2}(5\eta_{A}+6\eta_{\chi}-60)(\bar{\lambda}_{5}(N_{c}+4)-N_{c}(\bar{\lambda}_{4}+\bar{\lambda}_{3}))}{5N_{c}}$
$+\frac{1}{5}(\eta_{\chi}-5)\left(8\bar{\lambda}_{4}\bar{\lambda}_{5}-4\bar{\lambda}_{3}\bar{\lambda}_{5}(N_{\chi}-3)-\frac{\left(N_{c}^{2}(N_{\chi}+3)-2N_{c}(N_{\chi}-1)+16\right)}{N_{c}}\bar{\lambda}_{5}^{2}\right)$
$-\frac{1}{20}(\eta_{\psi}-5)\bar{\lambda}_{7}^{2}N_{\psi}\bigg]\,$, (27e)
$\partial_{t}\bar{\lambda}_{6}=$ $(2+\,\eta_{\psi}+\,\eta_{\chi})\bar{\lambda}_{6}$
$+\frac{1}{16\pi^{2}}\bigg[\frac{g^{4}\left(N_{c}^{2}-N_{c}-2\right)(10\eta_{A}+3(\eta_{\chi}+\eta_{\psi}-30))}{20N_{c}^{2}}$
$-\frac{g^{2}\bar{\lambda}_{7}\left(N_{c}^{2}-N_{c}-2\right)(5\eta_{A}+3(\eta_{\chi}+\eta_{\psi}-20))}{10N_{c}^{2}}$
$+\frac{3(\eta_{\chi}+\eta_{\psi}-10)}{10}\left(\bar{\lambda}_{6}^{2}+\frac{\left(N_{c}^{2}-N_{c}-2\right)}{2N_{c}^{2}}\bar{\lambda}_{7}^{2}\right)$
$-\frac{1}{5}(\eta_{\chi}-5)\Bigg(\bar{\lambda}_{4}\bar{\lambda}_{6}\left(N_{c}^{2}N_{\chi}-N_{c}N_{\chi}+2\right)+\bar{\lambda}_{3}\bar{\lambda}_{6}\left(N_{c}^{2}-N_{c}+2N_{\chi}\right)$
$\hskip 199.16928pt+\frac{2\left(N_{c}^{2}-N_{c}-2\right)}{N_{c}}\bar{\lambda}_{5}\bar{\lambda}_{6}\Bigg)$
$-\frac{2}{5}(\eta_{\psi}-5)\left(\bar{\lambda}_{1}\bar{\lambda}_{6}(N_{c}N_{\psi}+1)+\bar{\lambda}_{2}\bar{\lambda}_{6}(N_{c}+N_{\psi})\right)\bigg]\,$, (27f)
$\partial_{t}\bar{\lambda}_{7}=$ $(2+\,\eta_{\psi}+\,\eta_{\chi})\bar{\lambda}_{7}$
$+\frac{1}{16\pi^{2}}\bigg[\frac{g^{4}\left(N_{c}^{2}-4N_{c}-16\right)(10\eta_{A}+3(\eta_{\chi}+\eta_{\psi}-30))}{80N_{c}}$
@@
$+\frac{(\eta_{\chi}+\eta_{\psi}-10)}{5}\left(3\bar{\lambda}_{6}\bar{\lambda}_{7}-\frac{\left(N_{c}^{2}+3N_{c}+12\right)}{4N_{c}}\bar{\lambda}_{7}^{2}\right)$
$-\frac{2(\eta_{\psi}-5)}{5}\left(\bar{\lambda}_{1}\bar{\lambda}_{7}+\bar{\lambda}_{2}\bar{\lambda}_{7}N_{\psi}\right)$
$-\frac{2(\eta_{\chi}-5)}{5}\left(\bar{\lambda}_{4}\bar{\lambda}_{7}+\bar{\lambda}_{3}\bar{\lambda}_{7}N_{\chi}-\frac{\left(N_{c}^{2}(N_{\chi}+1)-2N_{c}(N_{\chi}+1)-4\right)}{10N_{c}}\bar{\lambda}_{5}\bar{\lambda}_{7}\right)\Bigg]\,.$ (27g)
@@
$\eta_{A}=$ $-\frac{\partial_{t}Z_{A}}{Z_{A}}=-\left.\frac{\partial_{p^{2}}\,{\rm tr}\left[{\cal P}^{(AA)}\partial_{t}\Gamma^{(AA)}_{k}\right]}{Z_{A}\,{\rm tr}\left[{\left({\cal P}^{(AA)}\right)}^{2}\right]}\right|_{p=0}=\frac{g^{2}}{16\pi^{2}}\bigg[\left(\frac{31}{36}\eta_{A}-\frac{11}{3}\right)N_{c}$
$\hskip 199.16928pt-\frac{(\eta_{\chi}-4)}{6}(N_{c}-2)N_{\chi}-\frac{(\eta_{\psi}-4)}{6}N_{\psi}\bigg]\,$, (28)
@@
and given that the anomalous dimensions enter in the regulated lines they lead to a closed system of equations which can be solved giving infinite-order expressions in a resumed form. This feature manifests one of the sources of the non-perturbative character of this fRG approach. The $\cal P^{(\ldots)}$ stand for the projectors onto the kinetic terms of the respective fields.
Furthermore, from the flows in (27), the previously discussed coefficients ($\bm{c}_{\ldots}$) can be read off as a function of the number of chiral matter. To facilitate the identification of each contribution, we have displayed the flows as a function of the fermion multiplicities given by the global symmetries,
@@
So far, we have not specified the pure gauge and Dirac parts of the action. In this work we employ a truncation in this sectors which includes the interaction terms up to dimension four and the classical tensor structures which is the same one as in Goertz:2024dnz . Furthermore, in the gauge fixing sector we have used the Landau gauge. As an additional cross-check, we have derived the flow of the gauge couplings from all the included avatars and obtained the correct one-loop structure.
@@
Figure 4: Value of the $\bm{c}_{{\rm A},i}$ coefficients which reflect the strength of the primer gauge corrections on the generation of the different four-fermion dressings. The sign crossing of $\bm{c}_{{\rm A},5}$ at $N_{c}\sim 11$ and $\bm{c}_{{\rm A},7}$ provides the identification of regimes and potential subdomains.
@@
Some further remarks are in order. First, note that only $\bm{c}_{{\rm C},ijk}$ depends on $N_{\textrm{gen}}$, because the only source of $N_{\textrm{gen}}$ dependence in the flow arises from pure fermionic loops, which contribute exclusively to $\bar{\lambda}_{i}\bar{\lambda}_{j}$ terms.
@@
Moreover, part of the system has a resemblance to QCD which permits to perform some cross checks. Operators ${\cal O}_{1}$ and ${\cal O}_{2}$ are equivalent to those known as (V-A) and (V-A)adj but here only with one chirality Gies:2005as ; Braun:2011pp ; Goertz:2024dnz . In fact, we find an agreement up to a factor 1/2 in the $N_{\psi}$ contributions in their respective flows, providing a consistency check to our computation.